### v1

# Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement

Abstract
LLMs are widely used for offensive language detection due to their advanced capability. However, the challenges posed by human annotation disagreement in real-world datasets remain underexplored. These disagreement samples are difficult to detect due to their ambiguous nature. Additionally, the confidence of LLMs in processing disagreement samples can provide valuable insights into their alignment with human annotators. To address this gap, we systematically evaluate the ability of LLMs to detect offensive language with annotation disagreement. We compare the binary accuracy of multiple LLMs across varying annotation agreement levels and analyze the relationship between LLM confidence and annotation agreement. Furthermore, we investigate the impact of disagreement samples on LLM decision-making during few-shot learning and instruction fine-tuning. Our findings highlight the challenges posed by disagreement samples and offer guidance for improving LLM-based offensive language detection.
Disclaimer: The samples presented by this paper may be considered offensive or vulgar.
Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement
Junyu Lu1, Kai Ma1, Kaichun Wang1, Kelaiti Xiao1 Roy Ka-Wei Lee2, Bo Xu1, Liang Yang1, Hongfei Lin1 1Dalian University of Technology, 2Singapore University of Technology and Design dutljy@mail.dlut.edu.cn, hflin@dlut.edu.cn
1 Introduction
Agreement Degree Example Hard Label Soft Label
$N^{++}$ Stand for something or else fall for anything. #BlackLivesMatter Hello world! What a great day to be alive #Trump2020 #MAGA 0 0
$N^{+}$ Come on man! Lock’em up!!! #maga Not the first time. You all misspelled #blacklivesmatter. Speak up! @user 0 0.2
$N^{0}$ Set fire to Fox News (metaphorically) @user is outing #BLACK-LIVES-MATTER as a cult! HE IS CORRECT! 0 0.4
$O^{0}$DISGUSTING #Democrats terrorize old folks just before #elections2020 I love this shit! #BlackLivesMatter 1 0.6
$O^{+}$user You’re a bumbling fool #elections2020 Elections 2020: Red Rapist v. Blue Racist 1 0.8
$O^{++}$ Y’all trending about kpop stans instead of #BlackLivesMatter big fack you Crazy idiots. This is batshit bullshit. #elections2020 1 1
Table 1: Eamples with different degrees of annotation agreement, where $N$ denotes non-offensive language and $O$ denotes offensive language. The superscripts ++, +, and 0 represent unanimous, medium, and low agreement, respectively. These samples are collected from the MD-Agreement dataset Leonardelli et al. (2021).
Offensive language on online platforms refers to socially considered inappropriate speech due to its vulgarity, obscenity, or irreverence Johnson and Lewis (2010), often inflicting psychological harm on individuals and society. As a result, the detection of offensive language has garnered significant attention from researchers in related fields Founta et al. (2018); Davidson et al. (2017).
Recently, researchers have started leveraging large language models (LLMs) to both detect and annotate offensive language, attracted by their extensive background knowledge and impressive reasoning capacity Kumar et al. (2024). Some studies Wang et al. (2023a); Huang et al. (2023) have shown that LLMs’ detection performance is comparable to that of humans. However, a critical gap in the current research is the idealization of offensive language classification into clear binary labels ("offensive" or "non-offensive"). This simplification overlooks the samples with annotation disagreement, which is common in real-world datasets Aroyo et al. (2019); Basile (2020); Uma et al. (2021b).
Compared to samples with unanimous agreement, those with annotation disagreement better reflect the complexity of human judgment in identifying offensive language Klenner et al. (2020); Leonardelli et al. (2021). These samples often feature ambiguous expressions, presenting challenges for both annotators and detector models Sandri et al. (2023). For example, in the low-agreement offensive sample ($O^{0}$) in Table 1, the statement links "Democrats" with "terrorizing old folks". While some annotators interpret it as inciting hostility due to its provocative tone, others view it as merely a strong emotional expression rather than offensive, leading to disagreement. This disagreement highlights the subjectivity and complexity involved in evaluating offensive language.
Furthermore, the decision confidence exhibited by LLMs when processing disagreement samples can offer valuable insights into their alignment with human annotators Weerasooriya et al. (2023); Baumler et al. (2023). An ideal detector should demonstrate higher confidence for samples with unanimous agreement, indicating clearer cases, and lower confidence for samples with low human agreement, which are inherently more ambiguous Uma et al. (2021a); Leonardelli et al. (2023). This alignment can further assess the feasibility of using LLMs to assist in offensive language detection and annotation tasks, revealing whether the model effectively captures the complexities and nuances that human annotators encounter.
In this paper, we explore the capabilities of LLMs in detecting offensive language with annotation disagreement. Specifically, we mainly focus on the following two research questions:
RQ1: How do LLMs perform in detecting offensive language with annotation disagreement?
How do disagreement samples affect LLMs during the learning phase?
For the first question, we evaluate the model in a zero-shot scenario from two perspectives: First, we assess the binary offensive language detection performance of various open-source and closed-source LLMs on samples with varying agreement degrees. Then, we analyze the relationship between annotation agreement and LLM confidence to gain deeper insights into the model’s decision-making process. For the second question, we focus on the impact of samples with varying agreement degrees during the learning phase, specifically examining two mainstream techniques: few-shot learning and instruction fine-tuning. We explore how samples with a single annotation agreement degree and different combinations of agreement categories affect the model’s detection performance. The main contributions of this paper are summarized as follows:
  • We provide a novel perspective on using LLMs for offensive language detection, emphasizing the importance of samples with annotation disagreement in research.
  • We evaluate the detection performance of multiple LLMs on disagreement samples, examining both binary accuracy and the relationship between annotators and LLMs.
  • We explore the impact of disagreement samples on LLM decision-making, with a focus on few-shot learning and instruction fine-tuning techniques.
2 Preliminary
2.1 Dataset
Given that disagreements may arise from both the inherent ambiguity of the data itself and low-quality annotations, the chosen benchmark dataset should meet two criteria: high annotation quality and the open availability of unaggregated annotation data. To this end, we use the high-quality offensive language corpus, MD-Agreement Leonardelli et al. (2021), as the dataset. This dataset consists of 10,753 samples collected from Twitter, with samples annotated by five trained annotators. In addition to the majority-voted hard labels, the dataset provides soft labels reflecting the annotation agreement. These soft labels are mainly categorized into three levels based on the degree of annotation agreement: unanimous agreement ($A^{++}$, i.e., agreement among all five annotators), mild agreement ($A^{+}$, i.e., four out of five annotators agreeing on the same label), and weak agreement ($A^{0}$, i.e., three annotators agreeing and two disagreeing). Moreover, offensive comments are categorized as $O^{++}/O^{+}/O^{0}$, while non-offensive comments are represented as $N^{++}/N^{+}/N^{0}$, respectively. Several samples with varying degrees of annotation agreement are provided in Table 1 for reference. The statistics of the MD-agreement dataset are shown in Table 2. The accuracy of its annotations has been validated by Sandri et al. (2023), with further details provided in Appendix A.1. This dataset is the official corpus of SemEval 2023 Task 11 Leonardelli et al. (2023) and has been widely utilized by researchers in the field Deng et al. (2023); Mokhberian et al. (2024).
Split $A^{++}$ $A^{+}$ $A$ $N^{++}$ $N^{+}$ $N$ $O$ $O^{+}$ $O^{++}$ N-Off. Off. Total
Train 2,778 1,930 1,884 2,303 1,295 1,032 852 635 475 4,630 1,962 6,592
Dev 464 317 322 346 199 171 151 118 118 1,103 387 1,103
Test 1,292 909 803 1,020 549 470 386 360 272 3,004 1,018 3,057
MD-Agreement 4,535 3,156 3,062 3,669 2,043 1,673 1,389 1,113 866 7,385 3,368 10,753
Table 2: Statistics of the MD-Agreement dataset, where $N$ denotes non-offensive language and $O$ denotes offensive language. The superscripts ++, +, and 0 represent unanimous, medium, and low agreement, respectively.
2.2 Models
To ensure a comprehensive evaluation, we access a range of LLMs, including both closed-source and open-source models: For closed-source models, we examine several widely used LLMs, including GPT-3.5, GPT-4, GPT-4o, GPT-o1111Since the current GPT-o1 API does not allow adjustment of the temperature coefficient, we only present its binary classification performance at a temperature of 1 in Table 3, which is the default setting. , Claude-3.5, and Gemini-1.5. For open-source models, we utilize three LLMs with different parameter sizes: LLaMa-3 8B and 70B, Qwen-2.5 7B and 72B, and Mixtral 8x7B and 8x22B. More information about the versions of LLMs is provided in Appendix A.4.
3 How Do LLMs Perform in Detecting Offensive Language with Annotation Disagreement?
In this section, we assess the performance of LLMs in a zero-shot scenario, focusing on both the binary accuracy in detecting offensive language and the relationship between annotation agreement and LLM confidence. We utilize all the samples from the MD-Agreement dataset for this evaluation and analyze the performance on subsets with different annotation agreements.
3.1 Evaluation of Binary Classification Performance
To evaluate the binary classification performance, we set the temperature coefficient of the LLM to 0, ensuring the model selects the category with the highest probability. We use accuracy and F1 score for assessment. The prompt template used for offensive language detection is listed in Appendix A.3 The outputs of the LLM are converted into hard predictions, with 1 indicating offensive and 0 indicating non-offensive. The results of LLMs in binary detection of offensive language are shown in Table 3. Based on the results, we can draw the following conclusions:
(1) Overall, in the zero-shot scenario without specific training, existing LLMs can fairly accurately identify samples with unanimous agreement ($A^{++}$). The average accuracy for closed-source and open-source LLMs is 88.28% and 86.07%, respectively. Furthermore, the detection performance of some outstanding open-source models, such as LLaMa3-70B, is now very close to that of widely used closed-source models. This trend underscores the increasing effectiveness of LLMs in detecting offensive language with clear labels.
(2) As the degree of annotation agreement decreases, the detection performance of LLMs consistently declines. For instance, GPT-4o achieves an F1 score of 85.24% for $A^{++}$ samples, but for mild agreement ($A^{+}$) and weak agreement ($A^{0}$) samples, the F1 scores drop by 10.64% and 28.18%, respectively. Notably, for weak agreement samples, the F1 scores of all LLMs fall below 65%. This finding highlights that samples prone to annotation disagreement pose significant challenges for LLMs in offensive language detection tasks.
(3) Larger model sizes generally improve the detection performance of LLMs, but the improvement is not as noticeable for samples with annotation disagreement. For example, in detecting $A^{++}$ samples, LLaMa3-70B outperforms LLaMa3-8B by 10.64%, but for $A^{0}$ samples, the improvement is only 3.11%. Similarly, for Mixtral and Qwen2.5, models with larger parameters (8$\times$22B and 72B) show nearly the same F1 scores as smaller models (8$\times$7B and 7B) on the $A^{+}$ and $A^{0}$ subsets. This suggests that the ability of the LLMs to recognize disagreement samples, which lie at the ambiguous boundary between offensive and non-offensive language, may not improve substantially with an increase in model size.
Overall $A^{++}$ $A^{+}$ $A^{o}$
Model Acc.\;$\uparrow$ F1\;$\uparrow$ Acc.\;$\uparrow$ F1\;$\uparrow$ Acc.\;$\uparrow$ F1\;$\uparrow$ Acc.\;$\uparrow$ F1\;$\uparrow$
Closed-Source Large Language Models  (CS-LLMs)
GPT-o1 78.35 69.03 91.95 81.29 77.50 72.03 59.08 58.63
GPT-4o 80.36 70.33 93.96 85.24 80.67 74.60 59.90 57.06
GPT-4 74.18 69.07 88.64 76.75 70.12 68.91 56.96 64.63
GPT-3.5 67.07 63.45 78.99 64.02 62.39 63.28 54.25 63.18
Claude-3.5 78.56 70.93 92.59 83.13 76.39 72.02 60.03 62.61
Gemini-1.5 69.50 66.07 83.53 69.70 64.48 65.73 53.89 64.07
Avg. of CS-LLM 74.67 68.15 88.28 76.69 71.93 69.43 57.35 61.70
Open-Source Large Language Models  (OS-LLMs)
LLaMa3-70B 76.93 71.06 91.40 81.36 74.46 71.96 58.03 64.37
LLaMa3-8B 71.82 65.31 85.56 70.72 68.22 66.06 55.19 61.26
Qwen2.5-72B 72.08 66.86 84.74 70.92 68.41 67.36 57.12 63.76
Qwen2.5-7B 71.10 67.14 85.34 72.02 66.92 67.25 54.31 64.06
Mixtral-8x22B 73.46 67.82 87.12 74.27 69.93 68.21 56.86 63.44
Mixtral-8x7B 70.57 65.59 82.27 67.58 67.14 66.32 56.76 63.63
Avg. of OS-LLM 72.66 67.30 86.07 72.81 69.18 67.86 56.38 63.42
Table 3: Binary classification performance of LLMs on the MD-Agreement dataset and its three subsets $A^{++}$, $A^{+}$, and $A^{0}$. Avg. of CS-LLM and OS-LLM respectively denote the average performance of the close-source and open-source LLMs. Results show the accuracy (Acc.) and $F_{1}$ in percentage (%). The bold and underline scores respectively represent the optimal and suboptimal values.
(4) Furthermore, we evaluate the accuracy of several LLMs for both offensive and non-offensive language with varying degrees of annotation agreement, as shown in Figure 2. The results indicate that for samples with the same degree of annotation agreement, most models demonstrate lower accuracy in identifying non-offensive language compared to offensive language. In particular, LLMs consistently exhibit the lowest detection accuracy for non-offensive samples with the lowest agreement ($N^{0}$), with an average accuracy of only 45.77%. This phenomenon can be attributed to the models’ tendency to classify these disagreement samples as the offensive language for safety reasons, reflecting an over-sensitivity of LLMs in the task of offensive language detection, a conclusion consistent with the findings of Kumar et al. (2024); Zhang et al. (2024). Furthermore, this behavior prevents LLMs from effectively capturing the ambiguous boundary between offensive and non-offensive language, much like human crowdsource annotators, further highlighting their limited capability in handling disagreement samples.
Figure 1: Accuracy of LLMs on detecting offensive and non-offensive language with different degrees of annotation agreement.
Overall $A^{++}$ $A^{+}$ $A^{o}$
Model MSE\;$\downarrow$ $\rho\uparrow$ MSE\;$\downarrow$ $\rho\uparrow$ MSE\;$\downarrow$ $\rho\uparrow$ MSE\;$\downarrow$ $\rho\uparrow$
Closed-Source Large Language Models (CS-LLMs)
GPT-4o 0.1138 0.6535 0.0514 0.8098 0.1268 0.6298 0.1928 0.2332
GPT-4 0.1716 0.6819 0.1131 0.7175 0.2064 0.5323 0.2224 0.2478
GPT-3.5 0.2163 0.5889 0.1878 0.6021 0.2430 0.4236 0.2309 0.1820
Claude-3.5 0.1306 0.6780 0.0657 0.7590 0.1544 0.5818 0.2022 0.2379
Gemini-1.5 0.2137 0.6305 0.1638 0.6970 0.2505 0.4517 0.2498 0.1877
Avg. of CS-LLMs 0.1692 0.6466 0.1164 0.7171 0.1962 0.5238 0.2196 0.2177
Open-Source Large Language Models (OS-LLMs)
LLaMa3-70B 0.1400 0.6990 0.0753 0.7634 0.1680 0.5856 0.2072 0.2369
LLaMa3-8B 0.1803 0.5912 0.1316 0.6533 0.2068 0.4572 0.2251 0.1667
Qwen2.5-72B 0.1909 0.6588 0.1380 0.6817 0.2235 0.5001 0.2359 0.2119
Qwen2.5-7B 0.1962 0.6024 0.1480 0.6638 0.2237 0.4756 0.2393 0.2056
Mixtral-8x22B 0.1810 0.6287 0.1251 0.6858 0.2112 0.4944 0.2326 0.2107
Mixtral-8x7B 0.1978 0.5921 0.1578 0.6267 0.2218 0.4709 0.2323 0.2132
Avg. of OS-LLMs 0.1810 0.6287 0.1293 0.6791 0.2092 0.4973 0.2287 0.2075
Table 4: Estimation of relationship between annotators and LLMs on MD-Agreement and its three subsets. Results show Mean Squared Error (MSE) and Spearman’s Rank Correlation Coefficient ($\rho$).
3.2 Evaluation of Relationship between Agreement Degree and LLM Confidence
We then evaluate the relationship between LLM confidence and human annotation agreement. The self-consistency method Chen and Mueller (2024); Wang et al. (2023b) is applied to measure the model’s confidence by resampling its outputs with varying temperature coefficients. To ensure a robust evaluation, we perform experiments across five temperature settings: 0, 0.25, 0.5, 0.75, and 1. The final soft predictions are obtained by averaging the hard predictions from the outputs. Mean Squared Error (MSE) is used as the metric to evaluate the alignment degree between LLM confidence and annotation agreement, following Uma et al. (2021a); Leonardelli et al. (2023), where a smaller MSE indicates closer alignment. We also use Spearman’s Rank Correlation Coefficient ($\rho$) to assess the statistical correlation between the two. The detailed introduction of metrics are provided in Appendix A.2. The results are presented in Table 4.
Figure 2: Self-consistency of several LLMs across varying degrees of annotation agreement with Cohen’s Kappa ($\kappa$) as the metric.
Based on the results, we observe that as annotation agreement decreases, the alignment degree between it and the LLM confidence shows a noticeable decline across multiple models. For example, GPT-4o, which demonstrates superior overall performance, has an MSE of about 0.05 for $A^{++}$ samples with unanimous agreement. However, when detecting $A^{0}$ samples with low agreement, the MSE increases to 0.2. The statistical correlation between the model’s confidence and annotation agreement also weakens accordingly. For $A^{++}$ samples, the correlation is generally high ($\rho>0.7$) for closed-source models, while for $A^{+}$ and $A^{0}$ samples with disagreements, the correlation is medium $(0.3<\rho\leq 0.7$) and low ($\rho<0.3$), respectively. These findings suggest that existing LLMs have not fully aligned with human standards for identifying offensive language, particularly for samples with annotation disagreement, where they struggle to capture the ambiguous nature.
Figure 3: Confusion matrix (raw counts and percentage) between soft predictions of GPT-4o (x-axis) and soft labels (y-axis).
We then analyze the self-consistency of the outputs generated by a single model across multiple samplings, aiming to gain deeper insights into its decision-making mechanism. We utilize Cohen’s Kappa ($\kappa$) as the metric, which is widely employed to assess the consistency of observations in classification tasks. The results of several LLMs’ self-consistency are shown in Figure 2. We note that, although the self-consistency of each model decreases as annotation agreement diminishes, the Kappa values between sampling remain above 0.75 even for $A^{0}$ samples, demonstrating good ($0.80.6\leq\kappa<$ 0.8) or even high ($\kappa\geq$ 0.8) consistency. This suggests that LLMs demonstrate high confidence in their offensive language detection decisions, but this can also lead to overconfidence, causing them to overlook the ambiguity of disagreement samples. This is particularly evident in the $A^{0}$ subset, where the consistency between LLMs and human annotators is notably low.
Furthermore, we construct a confusion matrix to analyze the relationship between the model’s soft predictions and the samples’ soft labels. GPT-4o is chosen as the representative model. The resulting confusion matrix is presented in Figure 3. The experimental results further reveal that even GPT-4, which achieves overall performance closest to the human level, tends to output the same prediction across multiple sampling attempts, regardless of the level of annotation agreement in the samples, exhibiting strong overconfidence in its predictions. This phenomenon underscores that relying on a single LLM lacks the diversity of judgment characteristic of human annotations in offensive language detection. Therefore, relying solely on LLMs to replace humans in detection and annotation remains a significant challenge.
Test Set $A^{++}$ $A^{+}$ $A^{o}$
Model Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$
GPT-4o (zero-shot) 80.11 0.1133 93.11 0.0560 79.76 0.1242 60.86 0.1923
[…]
$\qquad\qquad w/\;A^{+}$ 81.26 0.1144 93.65 0.0515 81.96 0.1157 61.80 0.2080
$\qquad\qquad w/\;A^{0}$ 81.22 0.1138 93.73 0.0558 82.18 0.1171 61.33 0.1979
[…]
$\qquad\qquad w/\;A^{++/0}$ 83.87 0.1063 95.74 0.0416 86.14 0.1079 63.55 0.2022
$\qquad\qquad w/\;A^{+/0}$ 82.07 0.1171 94.66 0.0500 82.18 0.1290 62.97 0.2059
$\qquad\qquad w/\;A^{++/+/0}$ 83.51 0.1045 95.51 0.0367 85.48 0.1078 63.32 0.2035
Table 5: Performance of GPT-4o on the test set of MD-Agreement in few-shot learning: Accuracy (Acc.) for binary classification and MSE for evaluating alignment degree between annotation agreement and model’s confidence. The first row shows GPT-4o’s performance in the zero-shot scenario, while the second and third sections evaluate the model with prompts containing a single level and combinations of agreement, respectively.
4 How Do Disagreement Samples Affect LLMs During the Learning Phase?
In this section, we examine the influence of samples with varying annotation agreement on the performance of LLMs during the learning phase, focusing on few-shot learning and instruction fine-tuning. We explore the effect of both single-level agreement samples and different agreement combinations on model performance.
4.1 Impact of Disagreement Samples on Few-shot Learning
We evaluate GPT-4o’s binary classification accuracy and the alignment between annotation agreement and the model’s confidence during the few-shot learning. To enhance the effectiveness of learning, we construct positive and negative sample pairs for each agreement degree, randomly selected from the MD-Agreement training set. Following the settings of Leonardelli et al. (2021), we assess the impact of both single-category agreement samples and combinations of samples with varying agreement degrees in the prompt, where the combined pairs use their respective single-category sample sets. For example, the marker $w/\;A^{++}$ refers to introducing sample pairs with unanimous agreement $A^{++}$ in the prompt, including one offensive and one non-offensive sample, while $w/\;A^{+/0}$ combines medium agreement $A^{+}$ and low agreement $A^{0}$ sample pairs, totaling four samples. The results for the test set of MD-Agreement and its three subsets with varying annotation agreement degrees are presented in Table 6. From these results, we draw the following conclusions:
(1) Overall, compared to the zero-shot scenario, few-shot learning can improve the model’s binary classification performance and its alignment with human annotations. Notably, the most explicit accuracy improvement, averaging 3.87%, is observed in the medium agreement subset ($A^{+}$), highlighting that few-shot learning enhances the model’s ability to better understand offensive language. However, on low agreement samples ($A^{0}$), introducing few-shot learning increases the MSE. This is because few-shot learning also boosts LLM confidence, leading to more consistent predictions, which do not align well with the more ambiguous human annotations in low-agreement samples.
(2) In the few-shot learning setup with a single annotation agreement category, using samples with annotation disagreement (i.e., $A^{+}$ and $A^{0}$) results in greater performance improvements across all subsets and evaluation metrics compared to using samples with unanimous agreement ($A^{++}$). This is because disagreement samples often represent the "edge cases” in the annotation process, which helps the model learn the decision boundaries that humans use to distinguish between offensive and non-offensive language. As a result, the model becomes more effective at understanding and detecting this complex linguistic phenomenon.
(3) Incorporating combinations with varying disagreement degrees into prompts can enhance model performance, though the effectiveness of different combinations varies. For GPT-4o, combining unanimous agreement samples with disagreement ones, e.g., $w/\;A^{++/0}$ and $w/\;A^{++/+}$, explicitly outperforms using only disagreement samples $w/\;A^{+/0}$. However, including combinations with multiple agreement degrees $w/\;A^{++/+/0}$ does not lead to further performance improvements; instead, the binary classification accuracy across various subsets is lower than that of $w/\;A^{++/0}$. This is because combining different types of samples increases the complexity of the model’s decision-making process, potentially hindering accurate judgments. Therefore, in few-shot learning, appropriately balancing samples with varying degrees of agreement is crucial for LLMs to effectively understand and detect offensive language.
To ensure the reliability of our conclusions, we replicate the experiment using the open-source LLM Qwen2.5-72B. The results show that the conclusions generally align with those of GPT-4o. Detailed results are provided in Appendix B.3.
Test Set $A^{++}$ $A^{+}$ $A^{o}$
Model Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$
LLaMa3-8B (zero-shot) 70.92 0.1856 85.22 0.1350 66.45 0.2167 54.09 0.2288
[…]
$\qquad\qquad w/\;A^{+}$ 77.04 0.1552 90.40 0.0898 74.81 0.1815 59.70 0.2291
$\qquad\qquad w/\;A^{0}$ 73.99 0.1348 86.53 0.1020 70.74 0.1537 56.31 0.1665
[…]
$\qquad\qquad w/\;A^{++/0}$ 79.29 0.1292 92.49 0.0641 78.11 0.1503 60.16 0.2138
$\qquad\qquad w/\;A^{+/0}$ 82.53 0.1075 95.20 0.0404 83.94 0.1160 61.80 0.1978
$\qquad\qquad w/\;A^{++/+/0}$ 84.23 0.1106 95.98 0.0379 85.81 0.1186 64.37 0.2150
Table 6: Performance of LLaMa3-8B on the test set of MD-Agreement under instruction fine-tuning.
4.2 Impact of Disagreement Samples on Instruction Fine-tuning
We select LLaMa3-7B as the backbone model and evaluate its performance after instruction fine-tuning. Specifically, following the settings of Leonardelli et al. (2021), we randomly sample an equal number of instances with different agreement levels from the training set of the MD-Agreement dataset for fine-tuning. We extract 1,800 samples each from the $A^{++}$, $A^{+}$, and $A^{0}$ categories, with this quantity based on the $A^{0}$ samples, as they are the least represented in the dataset. The instruction template remains consistent with that used in the zero-shot setting (see Appendix A.3). We also examine the impact of both single-category agreement samples and combinations of varying agreement degrees, adopting the same experimental markers as in Section 4.1. The results of the MD-Agreement test set are presented in Table 5. Based on these results, we can observe that:
(1) When fine-tuned with a single category of annotation agreement samples, LLaMA3-8B performs best with medium agreement $A^{+}$ samples, achieving strong classification accuracy and alignment with human annotations. In contrast, the effects of high-agreement $A^{++}$ samples and low-agreement $A^{0}$ samples explicitly differ: fine-tuning with $A^{++}$ enhances the model’s classification ability, while incorporating $A^{0}$ samples improves the alignment degree between LLM confidence and human annotation agreement, resulting in the lowest MSE. This phenomenon highlights the "double-edged sword” nature of disagreement samples in model detection performance. On the one hand, learning from these samples enables the model to capture the ambiguous boundary between offensive and non-offensive language while mitigating overconfidence in decision-making. On the other hand, relying solely on these challenging samples for fine-tuning may risk catastrophic forgetting, ultimately impairing overall performance. For this reason, $A^{+}$ samples provide the most optimal benefits, striking a well-balanced trade-off between effective learning and preventing model forgetting, enabling the model to enhance the ability without compromising overall performance due to excessive focus on complex samples.
(2) Fine-tuning with combinations of different agreement categories further improves the model’s detection performance compared to training with a single category of annotation agreement samples. Notably, training LLaMA3-8B with a mix of all three disagreement levels ($w/\;A^{++/+/0}$) achieves the best results, with overall performance comparable to GPT-4o in the few-shot learning (see Table 5). This underscores the advantage of training with a larger sample size, enabling a more comprehensive understanding and detection of offensive language. Additionally, among the combinations involving two categories, combining the two types of disagreement samples ($w/\;A^{+/0}$) shows explicitly better overall performance compared to the other two combinations, despite all of them using the same amount of instruction data. This further underscores the effectiveness of learning from disagreement samples and demonstrates that a well-balanced selection of different agreement levels during instruction fine-tuning can significantly benefit model performance.
We also replicate the instruction fine-tuning experiment with Qwen2.5-7B, using the same data for training and testing. The results confirm consistent conclusions, shown in Appendix B.4.
5 Related Work
Large Language Model. In recent years, large language models (LLMs) have rapidly emerged, showcasing extensive world knowledge and strong reasoning capabilities Kojima et al. (2022); Ouyang et al. (2022); OpenAI (2023). Many researchers have proposed diverse tasks to deeply analyze the relationship between the model’s outputs and human judgments Xu et al. (2024); Fan et al. (2024). In addition, the confidence of LLMs in their outputs has also attracted attention from researchers, which is often used to assess the reliability and robustness of the generated content Jiang et al. (2021). Various methods for estimating confidence have been proposed Zhang et al. (2020); Wang et al. (2023b); Tian et al. (2023); Lin et al. (2022). In this study, we employ the most straightforward approach, self-consistency, to estimate the model’s confidence.
Offensive Language Detection. Researchers have developed various methods for detecting offensive language Founta et al. (2018); Davidson et al. (2017); Mathew et al. (2021). As research advances, many studies argue that treating offensive language detection as a binary classification is an idealized assumption Basile et al. (2021); Basile (2020); Plank (2022), as annotation disagreement are inherent in datasets for such subjective task Pavlick and Kwiatkowski (2019); Uma et al. (2021b). Using majority voting for annotation agreement leads to information loss Davani et al. (2022), as these disagreements arise from the subtlety of the samples, not labeling errors Uma et al. (2022). Leonardelli et al. (2023) emphasizes that detection models should recognize this disagreement, rather than just improving classification performance.
Recently, several studies have begun evaluating the potential of LLMs for detecting offensive language Kumar et al. (2024); Roy et al. (2023), and designing detection methods based on them Park et al. (2024); Wen et al. (2023). Some studies Wang et al. (2023a); Huang et al. (2023) leverage the generative capabilities of LLMs to provide explanations for offensive language, assisting human annotation. Furthermore, Giorgi et al. (2024); Zhang et al. (2024) assess the sensitivity of LLMs to demographic information in the context of offensive language. Though great efforts have been made, these studies lack focus on the phenomenon of offensive language with annotation disagreement. In this paper, we aim to fill this research gap.
6 Conclusions
In this paper, we focus on the ability of LLMs to detect offensive language with annotation disagreement. We first evaluate the performance of multiple LLMs in a zero-shot scenario, examining both binary accuracy and the relationship between LLM confidence and annotation agreement. Compared to unanimous agreement samples, all LLMs exhibit explicitly lower detection accuracy for samples with annotation disagreement, failing to capture their ambiguous nature. Moreover, due to the inherent overconfidence, these models tend to make consistent decisions, resulting in a low degree of alignment with human annotators. These issues highlight the challenges posed by disagreement samples in offensive language detection and demonstrate that existing LLMs cannot fully replace humans in detection.
We further explore the impact of disagreement samples on LLM detection performance under few-shot learning and instruction tuning. Specifically, we analyze the effect of a single agreement degree and different combinations of agreement categories. Experimental results indicate that learning from disagreement samples notably improves detection accuracy and alignment degree with human annotations. This suggests that disagreement samples help the model learn the nuanced boundary between offensive and non-offensive language as defined by human society. Our study underscores the importance of disagreement samples in advancing offensive language detection with LLMs. We hope our findings and insights will contribute to the further development of related research.
Limitations
(1) Due to the scarcity of high-quality offensive language datasets with unaggregated labels, we only utilize the MD-Agreement dataset for experiments, which has been widely used in the field. Considering that relying on a single dataset may introduce bias or randomness, we mitigate this by conducting experiments with multiple closed-source and open-source LLMs to ensure the consistency and reliability of our findings, reducing the impact of bias. In future work, we plan to further explore the performance of LLMs in other subjective text analysis tasks, such as humor detection and misogyny detection, particularly in understanding samples with annotation disagreement.
(2) Due to usage restrictions, we are unable to evaluate the detection performance of several emerging LLMs, such as GPT-o3. We plan to further assess these more advanced models as soon as experimental conditions allow. Additionally, due to space limitations, the potential of certain techniques for detecting offensive language with annotation disagreement, such as reinforcement learning methods, are not discussed. We plan to explore these methods in future work and investigate effective strategies for enabling LLMs to fully leverage disagreement samples, thereby enhancing their detection capabilities.
(3) In evaluating the confidence of LLMs, we adopt a straightforward approach based on temperature resampling. We have noted another common method, the Logit-based approach Guo et al. (2017); Zhang et al. (2020), which involves using the logits of category-specific tokens to compute statistical probabilities within the model’s output. This method may provide deeper insights into the decision-making mechanisms of LLMs when handling disagreement samples. We plan to explore and evaluate this method in future work.
Ethics Statement
The opinions and findings contained in the samples of this paper should not be interpreted as representing the views expressed or implied by the authors. Accessing the MD-Agreement dataset requires users to agree to the creators’ usage agreements. The usage of these samples in this study fully complies with these agreements.
Appendix A Experimental Details
A.1 Details of Dataset
In this section, we provide a detailed introduction to the annotation quality control process of our used MD-Agreement dataset Leonardelli et al. (2021). The researchers implemented a two-stage annotation process: First, three linguists annotated a subset of the samples, and those with unanimous agreement were used as the gold standard for the annotation process. Following this, trained annotators from Amazon Mechanical Turk were employed to annotate the complete samples based on the established gold standard. After the task was completed, annotations from workers who did not achieve at least 70% accuracy were discarded. Additionally, it was ensured that each sample in the final dataset received five annotations. These measures help ensure the accuracy of the annotations. Sandri et al. (2023) further manually reviewed a random selection of 2,570 samples with annotation disagreement from the MD-Agreement dataset. The results showed that only 12 samples contained annotation errors, accounting for less than 0.5%, demonstrating the high quality and reliability of the dataset.
A.2 Description of Metrics
This section introduces the metrics used to assess the relationship between LLM confidence and the degree of human annotation agreement.
Mean Squared Error (MSE): The MSE is a widely used evaluation metric in regression tasks, measuring the difference between predicted and actual values. In this study, we adopt MSE for alignment estimation, as described by Leonardelli et al. (2023), where a smaller MSE indicates closer alignment between LLM confidence and agreement degree. We first obtain soft labels $y$ and soft predictions $\hat{y}$ of samples by averaging their discrete 0-1 annotation sequences $Y$ and the LLM outputs $\hat{Y}$ across different samplings, as follows:
$y_{i}=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\quad\hat{y_{i}}=\frac{1}{n}.\sum_{i=1}^{n}\hat{Y_{i}},$ (1)
where $n$ is the number of observations, set to $n=5$ in this paper, representing the number of annotators and LLM outputs. Then, the MSE is calculated as:
$MSE=\frac{1}{m}\sum_{i=1}^{m}(y_{i}-\hat{y}_{i})^{2},$ (2)
where $m$ is the total number of samples.
Range of Coefficient (($\rho$) Correlation Degree
$(0.7,1.0]$ High Correlation
$(0.3,0.7]$ Medium Correlation
$(0.0,0.3]$ Low Correlation
$0.0$ No Correlation
$[-1.0,0.0)$ Negative Correlation
Table A1: Correlation degree corresponding to different coefficient values ($\rho$).
Range of Kappa ($\kappa$) Agreement Degree
$[0.8,1.0]$ High Agreement
$[0.6,0.8)$ Good Agreement
$[0.4,0.6)$ Moderate Agreement
$(0.0,0.4)$ Poor Agreement
$0.0$ No Agreement
$[-1.0,0.0)$ Negative Correlation
Table A2: Agreement degree corresponding to different kappa values ($\kappa$).
Spearman’s Rank Correlation Coefficient ($\rho$): The Spearman’s Rank Correlation Coefficient is a non-parametric test that quantifies the degree of monotonic relationship between two variables. Unlike Pearson correlation, which assumes normally distributed variables, Spearman’s correlation does not require this assumption and can be applied to discrete data. This makes it an ideal choice for assessing the statistical correlation between annotation agreement and LLM confidence, which is computed as follows:
$\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)},$ (3)
where $di$ is the difference between the ranks of corresponding values of soft labels $y$ and predictions $\hat{y}$. The correlation degrees corresponding to different $\rho$ are shown in Table A1.
Cohen’s Kappa ($\kappa$): Cohen’s Kappa is widely used to evaluate the consistency between annotators, especially in classification tasks. Compared to simple percentage agreement, which directly calculates the proportion of agreement between two evaluators, Cohen’s Kappa provides a more precise measure as it accounts for the influence of random agreement. In this paper, we use Cohen’s Kappa to estimate the consistency of LLM outputs $\hat{Y}$. We first compute the kappa value between the $i$-th and the $j$-th LLM output of samples, i.e., $\hat{Y_{i}}$ and $\hat{Y_{j}}$, using the following formula:
$\kappa_{i,j}=\frac{P_{o_{i,j}}-P_{e_{i,j}}}{1-P_{e_{i,j}}},$ (4)
where $P_{o_{i,j}}$ represents the observed agreement, which is the proportion of agreement between $Y_{i}$ and $Y_{j}$, and $P_{e_{i,j}}$ refers to the expected agreement, calculated based on the probability of selecting categories, namely 0 or 1. We then calculate the average value of the kappas as follows, which is used as the metric:
$\kappa=\frac{1}{C(n,2)}\sum_{1\leq i<j\leq 5}\kappa_{i,j}$ (5)
The agreement degrees corresponding to different $\kappa$ are shown in Table A2.
Offensive language refers to words or expressions intended to insult, harm, or belittle individuals or groups. This includes:
$\cdot$ Insulting, derogatory, or combative language
$\cdot$ Discriminatory language based on race, religion, gender, or other personal attributes
$\cdot$ Profanity, violence, or vulgar expressions
$\cdot$ Sarcastic or harmful humor
Based on the above definition, determine if the following sentence is offensive language. Return only “yes” or “no”.
<If in the Few-shot Learning>
Example_1:
Input: [text of example_1]
Output: [label_1]
Example_2:
Input: [text of example_2]
Output: [label_2]
<Other Examples>
Here is the sample to be detected:
Input: [sample to be detected]
Output: [prediction]
Table A3: Prompt template of the LLM, consisting primarily of three parts: task definition, examples (only for the few-shot scenario), and the sample to be detected.
Model Version
GPT-o1 o1-preview-2024-09-12
GPT-4o gpt-4o-2024-08-06
GPT-4 gpt-4-turbo-2024-04-09
GPT-3.5 gpt-3.5-turbo-0125
Claude-3.5 claude-3-5-sonnet-20240620
Gemini-1.5 gemini-1.5-pro
[…]
LLaMa3-8B Meta-Llama-3-8B-Instruct
Qwen2.5-72B Qwen2.5-72B-Instruct
Qwen2.5-7B Qwen2.5-7B-Instruct
Mixtral-8x22B Mixtral-8x22B-Instruct-v0.1
Mixtral-8x7B Mixtral-8x7B-Instruct-v0.1
Table A4: Specific versions of used LLMs.
A.3 Design of Prompt Template
To enhance the reproducibility of our study, we avoided conducting complex prompt engineering. Instead, we directly referenced Roy et al. (2023) to design a straightforward prompt template, as shown in Table A3. The template includes three parts: first, the definition of offensive language, which aligns with that used in the MD-Agreement dataset Leonardelli et al. (2021) to ensure the accuracy of the evaluation; second, examples of varying degrees of disagreement in a few-shot scenario; and finally, the sample to be detected.
A.4 Other Experimental Settings
We access closed-source LLMs via their official APIs and deploy open-source LLMs with parameters downloaded from Hugging Face  [^2] https://huggingface.co/. To ensure a fair comparison, we use model versions released around the same time, as detailed in Table A4. Except for the temperature coefficient, other hyperparameters, such as top-p and top-k, are set to their default values for each model. For instruction fine-tuning, we adopt the efficient Qlora fine-tuning method. The learning rate is set to 2e-4, with a per-device batch size of 36. We train the model for 15 epochs using the AdamW optimizer, applying an early stopping mechanism. We reserve the parameters of best-performing models based on the development set and evaluate their performance on the test set. The models are trained on two NVIDIA H100 80GB GPUs. All the few-shot learning and instruction fine-tuning experiments are repeated five times with different random seeds to minimize error, and the average results are reported.
A.5 Handling of Refusal Behavior
Handling offensive language can trigger the refusal behavior of LLMs, as they are designed with ethical and safety considerations Kumar et al. (2024). Nevertheless, in our experiments, refusal occurred only in the zero-shot evaluation setting, where Claude-3.5, with a temperature coefficient set to 1, failed to generate responses for 23 samples. When the experiment was repeated with the same settings, the model successfully provided predictions for these samples. This phenomenon also highlights the model’s sensitivity to offensive language and the non-deterministic nature of its outputs.
Appendix B Supplementary Experiments
Figure B1: Accuracy of LLMs on detecting offensive language with different degrees of annotation agreement under different temperature sampling settings.
Figure B2: Consistency of outputs from different LLMs across varying degrees of annotation agreement with Cohen’s Kappa as the metric. The color scale represents different Kappa values.
B.1 Impact of Temperature Sampling on Detection Performance of LLMs
In this section, we analyze the impact of temperature sampling on the accuracy of detecting offensive language by LLMs. We select four representative models for comparison: the closed-source models GPT-4 and Claude-3.5, as well as the open-source models LLaMa3-70B and Qwen2.5-72B. The experimental results are shown in Figure B1. Based on these results, we conclude that after adjusting the temperature coefficient, the detection accuracy of each LLM remains generally stable, although some fluctuations are observed, with varying degrees of sensitivity to the temperature coefficient across different models. As the temperature increases, the accuracy of most models shows a declining trend, with the sole exception being Qwen2.5-72B, which exhibits an increase in accuracy. This may be due to differences in the models’ training mechanisms. Nevertheless, the performance ranking between the models remains stable, indicating that changes in the temperature coefficient do not notably affect the performance differences among the models.
B.2 Consistency Analysis Across Different LLMs
Building upon Section 3.2, we further explore the consistency of hard predictions across different LLMs when processing samples with varying degrees of annotation agreement. We select six representative models, including the close-source models GPT-4o, Claude-3.5, and Gemini-1.5, as well as the open-source models Mixtral-8x22B, LLaMa3-70B, and Qwen2.5-72B. Cohen’s Kappa is used as the metric. The results are presented in Figure B2. Based on the results, we can observe that:
As annotation agreement decreases, cross-model consistency in detecting offensive language declines more significantly compared to each model’s self-consistency. For unanimous agreement samples ($A^{++}$), cross-model consistency generally exhibits good agreement, with $\kappa>0.6$. However, for low agreement samples $A^{0}$, consistency drops explicitly, with many models showing poor agreement ($\kappa<0.4$), despite many of these models exhibiting similar overall performance in terms of both binary classification accuracy and alignment with human annotations (see Table 3 and 4). Notably, the lowest prediction consistency Kappa is only 0.28 between GPT-4o and Gemini 1.5. A potential reason for this phenomenon is that different models are trained on diverse datasets and undergo distinct value alignment processes, resulting in varying sensitivity to contextual features. In future work, we will explore the relationship between cross-model consistency and human annotation agreement in offensive language detection. Additionally, we aim to investigate the potential of leveraging multiple LLMs for collaborative annotation of offensive language.
Overall $A^{++}$ $A^{+}$ $A^{o}$
Model Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$
Qwen2.5-72B (zero-shot) 72.08 0.1962 84.74 0.1480 68.41 0.2237 57.12 0.2393
[…]
$\qquad\qquad w/\;A^{+}$ 79.10 0.1275 91.95 0.0702 78.66 0.1414 60.16 0.1993
$\qquad\qquad w/\;A^{0}$ 82.56 0.1054 94.12 0.0514 83.50 0.1088 64.14 0.1832
[…]
$\qquad\qquad w/\;A^{++/0}$ 82.43 0.1099 93.58 0.0530 84.38 0.1108 63.55 0.1950
$\qquad\qquad w/\;A^{+/0}$ 82.96 0.1044 94.97 0.0427 84.71 0.1090 62.97 0.1927
$\qquad\qquad w/\;A^{++/+/0}$ 82.04 0.1101 93.42 0.0544 84.05 0.1095 62.73 0.1949
Table B1: Performance of Qwen2.5-72B on the test set of MD-Agreement in few-shot learning.
Overall $A^{++}$ $A^{+}$ $A^{o}$
Model Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ MSE\;$\downarrow$
Qwen2.5-7B (zero-shot) 69.77 0.1998 83.44 0.1542 66.12 0.2289 53.04 0.2379
[…]
$\qquad\qquad w/\;A^{+}$ 80.41 0.1347 93.11 0.0649 80.64 0.1497 62.73 0.2211
$\qquad\qquad w/\;A^{0}$ 80.08 0.1395 92.57 0.0704 80.09 0.1533 60.86 0.2238
[…]
$\qquad\qquad w/\;A^{++/0}$ 82.17 0.1192 94.50 0.0486 84.27 0.1243 63.90 0.2082
$\qquad\qquad w/\;A^{+/0}$ 81.88 0.1185 94.43 0.0522 83.39 0.1283 63.90 0.2060
$\qquad\qquad w/\;A^{++/+/0}$ 83.91 0.1149 95.82 0.0395 85.59 0.1209 65.42 0.2181
Table B2: Performance of Qwen2.5-7B on the test set of MD-Agreement under instruction fine-tuning.
B.3 Few-shot Learning with Qwen2.5-72B
We replicate the few-shot learning experiment from Section 4.1 using the open-source LLM Qwen2.5-72B, employing the same sample pairs in the prompts. The results are shown in Table B1. Based on the results, we observe the following:
In the few-shot learning with samples of varying annotation agreement degrees, the results of Qwen2.5-72B align closely with the trends of GPT-4o (see Table 5). Whether introducing samples with a single annotation agreement degree or combinations of different agreement categories, the detection performance of the model shows notable improvement compared to the zero-shot scenario. Additionally, the benefit to model performance varies explicitly depending on the annotation agreement degree and the combinations used as prompts.
Furthermore, compared to GPT-4o, Qwen2.5-72B demonstrates two distinct differences: (1) On the subset of low-agreement samples, the introduction of few-shot learning also results in a noticeable improvement in the alignment between LLM confidence and annotation agreement. This is primarily because Qwen2.5-72B performs less effectively than GPT-4o in detecting offensive language under zero-shot scenarios, making it more susceptible to performance enhancements through few-shot learning. (2) For Qwen2.5-72B, the combination of disagreement samples, i.e., $w/\;A^{+/0}$, achieves the best performance across most subsets and metrics. In contrast, GPT-4o performs better with combinations such as $w/\;A^{++/+}$ and $w/\;A^{++/0}$. This highlights that the effects of learning from disagreement samples differ between different LLMs, which is closely related to the ambiguous characteristics of these samples.
B.4 Fine-tuning with Qwen2.5-72B
We replicate the instruction fine-tuning experiment from Section 4.2 using Qwen2.5-7B, training with the same instruction data. Based on the results shown in Table B2, we observe conclusions that are largely consistent with those on LLaMa3-8B. Specifically, Qwen2.5-7B performs best with medium agreement ($A^{+}$) when fine-tuned using a single category of annotation agreement across various subsets and metrics. Compared to unanimous agreement samples ($A^{++}$), training with disagreement samples helps the model’s confidence better align with the degree of human annotation agreement, resulting in a lower MSE. When fine-tuning with combinations of different agreement categories, using lower agreement combinations (e.g., $w/\;A^{++/0}$ and $w/\;A^{+/0}$) leads to higher detection accuracy and closer alignment with the annotations, compared to higher agreement combinations ($w/\;A^{++/+}$), especially in the detection of disagreement subsets. These findings further demonstrate the importance of learning from disagreement samples during instruction fine-tuning, which enhances the model’s understanding and judgment of offensive language, particularly for ambiguous samples.
B.5 Error Analysis
Exp. Sentence Label GPT4o LLaMa3
(a) Why does the Minneapolis mayor look like a more Soy version of Justin Trudeau $O^{+}$ $O^{0}$ $N^{0}$
(b) This is Amerikkka! #blacklivesmatter <url> $N^{0}$ $O^{++}$ $O^{++}$
(c) Hey #BlackLivesMatter SAY THEIR NAMES! Look these are YOUR PEOPLE! Say Their Names! Do It! Where is Your Outrage Now?! Come forth You #Hypocrites! Did They Die In Vain?! Ye Generation Of Vipers! These Are Your People Murdered! $O^{0}$ $N^{0}$ $N^{++}$
(d) Doesn’t have a kid’s cancer charity he can draw from? #Elections2020 $N^{+}$ $O^{++}$ $N^{++}$
Table B3: Error analysis of samples with annotation disagreement in offensive language detection.
To gain deeper insight into the challenge posed by offensive language with annotation disagreement, we manually inspect the set of samples misclassified by the models. The following two main types of errors are identified, with samples and predictions from GPT-4o and LLaMa3-72B shown in Table B3 for illustration:
Type I error refers to samples that are labeled as non-offensive but are detected as offensive. This error primarily arises from subtle linguistic features such as sarcasm and metaphor, which make the judgment of the sample ambiguous. For instance, in Example (a), the term “Amerikkka” is a variant of “America” used to intensify emotional expression. Due to insufficient context, most annotators do not consider it offensive. However, GPT-4o and LLaMa3, due to their sensitivity to the hashtag #blacklivesmatter, consistently classify it as offensive language. Similarly, in Example (b), a sarcastic rhetorical question leads to a misclassification by GPT-4o. This phenomenon highlights the complexity that human annotators face in determining offensive language and also reveals the issue of over-sensitivity in existing LLMs to certain linguistic expressions, resulting in decisions that do not align with human standards. In future work, we will perform a more detailed analysis of expressions in samples with disagreement annotation and explore how different types of expressions affect model detection performance.
Type II error refers to sentences labeled as offensive but classified as non-offensive by the models. This error primarily arises from the models lacking or failing to effectively integrate the necessary background knowledge for detecting offensive content, leading to an inaccurate understanding of the sample’s true meaning. For example, in Example (c), the comparison between the mayor of Minneapolis and Justin Trudeau uses “Soy” as an adjective, which implies weakness and is intended to belittle the mayor. Both human annotators and GPT-4o capture the offensive nature of the sample, but LLaMa3 fails to correctly identify its offensiveness due to insufficient relevant knowledge. In Example (d), the phrase “Ye Generation of Vipers”, a religiously charged expression, is used to strongly criticize police brutality against black people. However, the model fails to integrate the context, leading to a missed detection. We plan to introduce more comprehensive background knowledge to enhance the understanding capability of LLMs and explore the performance of knowledge-enhanced models in detecting disagreement samples.